OpenStudy (anonymous):

Can someone please help me! I know I'm messing something up but I'm not sure where! The question is: Find a particular solution of the differential equation (-5/4)y''−2y'+y=2xe^(4x) using the Method of Undetermined Coefficients (primes indicate derivatives with respect to x).

OpenStudy (anonymous):

i found r to be -2 and(2/5) hihc i then found the complementary solution to be yc=c1e^(2/5x)+c2e^-2x yp=Ae^x +Be^x

OpenStudy (anonymous):

y'p=the same thing as yp and so does y''p(x)

OpenStudy (anonymous):

then i plugged it in and ended up with e^x ((-9/4)A-(9/4)B)=2xe^x

OpenStudy (anonymous):

which i assume is wrong because I cannot solve for the variable

OpenStudy (phi):

shouldn't it be \[ y_p=Axe^{4x} +Be^{4x} \]

OpenStudy (anonymous):

omg yes! thanks!